Please help me solve this problem.

4a+9b=8
8a+6c=-1
6b+6c=-1

thanks for your help!!

Answer by stanbon(75887)

4a + 9b + 0 = 8
8a + 0 + 6c= -1
0 + 6b + 6c= -1
------------
Subtract twice the 1st equation from the 2nd to get:
4a + 9b + 0 = 8
0 - 18b + 6c= -17
0 + 6b + 6c= -1
---------------
Subtract the 3rd equation from the 2nd to get:
4a + 9b + 0 = 8
0 - 24b + 0= -16
0 + 6b + 6c= -1
------------------------
Solve for "b" in the 2nd equation:
b = 16/24 = 2/3
--------------------
Substitute into the 1st to solve for "a" and into the 3rd to solve for "x".
4a + 9(2/3) = 8
4a + 6 = 8
4a = 2
a = 1/2
-------------
6b + 6c = -1
6(2/3) + 6c = -1
4 + 6c = -1
6c = -5
c = -5/6
===================
a = 1/2
b = 2/3
c = -5/6
===================
Cheers,
Stan H